### v1

# SEDMamba: Enhancing Selective State Space Modelling with Bottleneck Mechanism and Fine-to-Coarse Temporal Fusion for Efficient Error Detection in Robot-Assisted Surgery

Abstract
Automated detection of surgical errors can improve robotic-assisted surgery. Despite promising progress, existing methods still face challenges in capturing rich temporal context to establish long-term dependencies while maintaining computational efficiency. In this paper, we propose a novel hierarchical model named SEDMamba, which incorporates the selective state space model (SSM) into surgical error detection, facilitating efficient long sequence modelling with linear complexity. SEDMamba enhances selective SSM with bottleneck mechanism and fine-to-coarse temporal fusion (FCTF) to detect and temporally localize surgical errors in long videos. The bottleneck mechanism compresses and restores features within their spatial dimension, thereby reducing computational complexity. FCTF utilizes multiple dilated 1D convolutional layers to merge temporal information across diverse scale ranges, accommodating errors of varying durations. Besides, we deploy an established observational clinical human reliability assessment tool (OCHRA) to annotate the errors of suturing tasks in an open-source radical prostatectomy dataset (SAR-RARP50), constructing the first frame-level in-vivo surgical error detection dataset to support error detection in real-world scenarios. Experimental results demonstrate that our SEDMamba outperforms state-of-the-art methods with at least 1.82% AUC and 3.80% AP performance gain with significantly reduced computational complexity. [^1] The corresponding dataset, code and models will be released upon paper acceptance.
Index Terms:
surgical error detection, robotic-assisted surgery, selective state space model, video, temporal
I Introduction
Robotic-assisted surgery (RAS) is well-adopted globally and the preferred option for many surgical specialties (e.g. urology) [1]. RAS is complex, highly variable and requires advanced technical and cognitive skills. Adverse events during RAS and broadly laparoscopic surgery are not uncommon [2]. A proportion of these are operator’s errors leading to consequences ranging from suboptimal execution of surgical tasks to patient injury and rarely death. Overall, human errors are the second cause of injury and death in RAS [3].
Observational clinical human reliability assessment (OCHRA) is a standardized method to assess the quality of surgical execution by detecting and characterising the nature (procedural or executional) and severity of technical errors on various specialties [4, 5, 6]. Clinical validation has identified executional errors, following the OCRA definitions, as a strong predictor of both skill and patient outcomes [7, 8]. This strongly promotes the application of such tools for reducing adverse events and surgical training. Deployment of OCHRA is impractical due to the significant time required for reviewing and manual annotation. Subjectivity in validity and reliability issues further prioritize the need for automated solutions [9, 10].
Propelled by deep learning (DL) methods and the public JIGSAWS dataset [11], containing atomic gesture and skill annotations from RAS tasks in dry labs, significant prior work on RAS video understanding for skill analysis focused on recognition and temporal localization of surgical gestures [12, 13, 14, 15, 16]. Recent work adapted DL methods for gesture recognition to in-vivo datasets[17, 18, 19], and also studied the association of the observed gesture sequence to patient outcomes and surgeon’s skill [20]. Detection and temporal localization of procedural and executional errors have not been extensively studied. Hutchinson et al. [21] incorporated error annotations in two JIGSAWS tasks (simulation suturing and needle passing). Procedural errors are considered when a surgical gesture is omitted or the task is carried out with a gesture sequence that deviates from the defined grammar graph, while executional errors denote moments of suboptimal performance. For each gesture, a maximum of four errors were considered from a list (multiple attempts, needle orientation, needle drop, out of view), used to characterize the entire (begin to end) gesture as erroneous (if any error was noted) or normal [21]. Employing these annotations, Li et. al. compared three DL classifiers for detecting erroneous gestures [22].
We argue that the annotation approach in [21], though successful in dry-lab, is not directly applicable to in-vivo RAS settings. It is not possible to consider the same error definitions for all gestures. It is also typical for errors to span across multiple gestures, thus their temporal localization can not always align to the temporal boundaries (begin, end) of gestures. Consequently, its annotation based on pre-segmented gesture clips fails to accurately provide temporal boundaries for the start and end of errors. In this work, we define custom error descriptions, following OCHRA, and present the first in-vivo RAS dataset with frame-level error annotations to provide more accurate temporal boundaries of errors for surgical error detection (SED).
SED aims to detect errors that occur during surgery. In [22], kinematic data was further downsampled to lengths of 30 using sliding windows within manually pre-segmented gesture clips, serving as the input for a CNN-LSTM network to detect executional errors. However, its parallel computing capability is constrained by long short-term memory (LSTM). Samuel and Cuzzolin [23] developed a deep convolutional autoencoder to detect errors in pre-trimmed video clips in an unsupervised manner, yet the local receptive field of convolution layers hinders its global temporal modelling capability. Overall, existing research on SED has primarily focused on short kinematics data and video clips, efficient SED in long surgical videos remains unexplored. To achieve this objective, it is essential to consider two key factors: long-term dependency modelling and computational complexity. Transformer architectures, benefiting from the self-attention mechanism, enable global modelling and have demonstrated impressive performance in surgical video analysis tasks such as surgical phase recognition [24, 25]. Nevertheless, the quadratic computational complexity of self-attention with respect to video sequence length significantly increases the computational burden, impeding its efficiency in handling long video scenarios.
Figure 1: A surgical video from the SAR-RARP50 dataset, including errors of varying durations.
Recently, state space models (SSMs) have proven highly effective in capturing long sequences in the natural language processing domain [26, 27]. Selective SSM [27], namely Mamba, distinguishes itself from other SSM-based methods by integrating a data-dependent SSM with a selection mechanism that employs parallel scanning, thereby achieving linear complexity while preserving global receptive fields. Inspired by the great success of Mamba, researchers have explored it in the computer vision domain, yielding promising performance. Zhu et al. [28] proposed a bidirectional Mamba (Vim) that scans images forward and backward to enhance spatial-aware understanding for image tasks, significantly improving computational efficiency while achieving higher performance compared to well-established Transformers. SegMamba [29] constructed a tri-oriented Mamba for 3D medical image segmentation. However, as depicted in Fig. 1, surgical videos contain errors with varied durations, necessitating the algorithm to consider temporal information at different granularities. Prior selective SSM-based works rarely consider modelling multi-scale temporal information, we thus showcase the effectiveness of a novel selective SSM-based method named SEDMamba, which captures temporal information across various scales to detect surgical errors in suturing from robotic-assisted radical prostatectomy (RARP) cases.
This work makes the following contributions:
  1. 1.
    We propose a novel hierarchical model, SEDMamba, constructed with bottleneck multi-scale state space blocks to detect surgical errors in long videos. To handle errors with variable durations, we design a fine-to-coarse temporal fusion (FCTF) using different dilated 1D convolutional layers to merge temporal information across different scale ranges. Besides, to further reduce computational complexity, we incorporate a bottleneck mechanism that compresses the spatial dimension of features, enabling selective SSM to concentrate on modelling the temporal dimension. To the best of our knowledge, this is the first attempt to exploit selective SSM in surgical error detection.
  2. 2.
    We present the first frame-level surgical error annotations of an in-vivo RAS dataset.
  3. 3.
    The proposed SEDMamba outperforms state-of-the-art video analysis methods and selective SSM-based methods by 1.82-13.06% AUC and 3.8-18.86% AP while requiring 11.71-75.79% fewer model parameters and 40.71-90.93% fewer FLOPs.
II Methodology
II-A Dataset Description and Annotation
A total of 48 videos with 60 Hz from the open-source, SAR-RARP50 [30, 31], dataset of the RARP cases were independently and blindly reviewed by two assessors. SAR-RARP50 focuses on the suturing of the dorsal venous complex (DVC) task and contains open-source annotations of eight gestures used as a task analysis sheet [30]. We developed a customized tool for the annotation of both procedural and executional errors using a previously validated suturing checklist and the OCHRA methodology [32]. This was summarized as a list of possible procedural and executional errors within any given suturing task, as seen in Table I. The duration of these videos varies from one minute to over eleven minutes, including the complete suturing process. Two assessors (one senior and one junior RAS surgeon with expertise in RAS education) analyzed these untrimmed videos sequentially frame by frame and annotated the start and end frame indexes of all errors. Fig. 2 illustrates three types of errors from our annotations. Inter-rater reliability was verified with Pearson’s correlation showing very strong agreement ($r$=0.8726, $p<$0.001) between the two assessors. Finally, consensus was established for all disagreements via joint review.
Figure 2: Example error annotations at SAR-RARP50: (a) E6-Incorrect angle grasping needle; (b) E5-Tissue damage; (c) E24-Poor instrument control (clashing).
TABLE I: Error coding descriptions used for annotation
Error Type Error Description Incidence Total Frames
E1
Multiple attempts (end of 1st attempt
to end of successful attempt) i.e. multiple piercings,
not regrasping with needle held in tissue
510 119635
E2 Needle drop/slip if in tissue 77 3673
E3 Instrument(s) out of view 1 27
E4 Needle out of view 259 32548
E5 Tissue damage including poor/error in tissue stabilisation 113 16799
E6 Incorrect angle grasping needle (not perpendicular) 116 2237
E7 Incorrect position along needle (normally 2/3rds) 89 1737
E8 Excessive force 65 6220
E9 Needle does not follow the curve 169 15078
E10 Needle entry incorrect angle i.e. not perpendicular 35 2761
E11 Grasped at needle tip 106 888
E12 Suture is loosened 0 0
E13 Thread caught in instrument 67 14423
E14 Knot tied is not square (C/Reverse C) 0 0
E15 Inadequate number of throws 1 93
E16 Suture pulled through tissue before tying knot 0 0
E17 Incorrect distancing between needle drives (too close/far) 4 977
E18 Suture not pulled through between needle drives 4 886
E19 Suture entanglement 34 20364
E20 Fraying the suture 5 600
E21 Snapping the suture 4 67
E22 Dangerous/poor/incorrect needle disposal 0 0
E23 Incorrect/poor camera control (blurred/poor view) 162 34204
E24
Incorrect/poor instrument control (clashing,
3rd arm/non-dominant use)
686 47138
A total of 2507 error cases were annotated, with E24-“Incorrect/poor instrument control” occurring the most frequently (686), followed by E1-“Multiple attempts” (510). The frequency of error types had statistically significant difference ($\chi 2$ = 1945.15, $p<$ 2.2e-16). The duration of errors varies, with the longest average duration being E19-“Suture entanglement”, averaging around 599 frames per case, while the shortest is E11-“Grasped at needle tip”, averaging 8 frames per case. The 48 videos comprise a total of 976037 frames. In the SED task, the frame is categorized as “normal” if no error exists or “error” if any type of error exists, resulting in 320355 error frames. To detect errors of varying durations, irrespective of their types, it is particularly crucial to model temporal information across multiple scales.
II-B Surgical Error Detection Model (SEDMamba)
This section introduces the preliminary concepts related to SEDMamba, including state space modelling and discretization. We then formulate the SEDMamba architecture and its core bottleneck multi-scale state space (BMSS) block with fine-to-coarse temporal fusion.
Figure 3: The overall architecture of the proposed SEDMamba. (a) The fundamental block of SEDMamba, namely the bottleneck multi-scale state space (BMSS) block; (b) Fine-to-coarse temporal fusion (FCTF); (c) Bottleneck mechanism.
II-B1 Preliminaries
State space models (SSMs), such as structured state space sequence models [26] and selective SSM [27], are inspired by continuous system that map 1D feature or sequence $x(t)\in\mathbb{R}^{L}$ with length $L$ to $y(t)\in\mathbb{R}^{L}$ through a hidden state $h(t)\in\mathbb{R}^{N}$. Formally, SSMs can be formulated as:
$h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t);y(t)=\mathbf{C}h(t)$ (1)
where $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the evolution parameter, $\mathbf{B},\mathbf{C}\in\mathbb{R}^{N\times 1}$ denote as the projection parameters, $N$ is the state size. To transform the continuous parameters $\mathbf{A},\mathbf{B}$ in Eq. 1 to discrete $\overline{\mathbf{A}},\overline{\mathbf{B}}$, SSMs generally adopt the zero-order hold rule including a timescale parameter $\Delta$ for discretization:
$\overline{\mathbf{A}}=\exp(\Delta\mathbf{A}),\overline{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A})-\mathbf{I})\cdot\Delta\mathbf{B}$ (2)
Eq. 1 can now be rewritten as a discretized version based on Eq. 2:
$h_{t}=\overline{\mathbf{A}}h_{t-1}+\overline{\mathbf{B}}x_{t};y_{t}=\mathbf{C}h_{t}$ (3)
The output $y$ of the model can be computed through a global convolution for efficient training:
$\overline{\mathbf{K}}=\left(\mathbf{C}\overline{\mathbf{B}},\mathbf{C}\overline{\mathbf{AB}},\ldots,\mathbf{C}\overline{\mathbf{A}}^{L}\overline{\mathbf{B}},\ldots\right);y=x*\overline{\mathbf{K}}$ (4)
where $L$ is the length of the input sequence $x$, and $\overline{\mathrm{K}}\in\mathbb{R}^{L}$ is a structured convolutional kernel [26].
II-B2 Overview of SEDMamba
SEDMamba has a hierarchical architecture, shown in Fig. 3, comprising three bottleneck multi-scale state space (BMSS) blocks. Taking a video $V\in\mathbb{R}^{L\times C\times H\times W}$ with $L$ length, we first extract the spatial embedding sequence $\mathcal{E}\in\mathbb{R}^{L\times D}$ from a fixed DINOv2 [33], where $D$ is the spatial dimension. DINOv2 is a strong vision foundation model proven to extract task-agnostic visual spatial embedding and excel in video-based tasks [33]. Then, SEDMamba compresses $\mathcal{E}$ in the spatial dimension and captures fine-to-coarse temporal feature through each BMSS block, resulting in $i$-th hierarchical spatial-temporal representations with size $L\times\frac{D}{2^{i}}$. The 1D convolutional layer with 1 kernel size and 1 channel is then used for binary classification and outputs the final error probability prediction $P\in\mathbb{R}^{L\times 1}$. Note that each value in $P$ indicates the likelihood that the corresponding frame is erroneous. We use binary cross-entropy loss $\mathcal{L}_{bce}$ for supervised training.
II-B3 Bottleneck Multi-scale State Space (BMSS) Block
The structure of the BMSS block is shown in Fig. 3(a). Encouraged by the strong long-range modelling ability of SSMs, the BMSS block incorporates selective SSM with bottleneck mechanism and fine-to-coarse temporal fusion (FCTF) to reduce computational complexity and capture inter-frame long-range dependencies across multiple temporal scales based on spatial embedding sequences, respectively. Given input $x\in\mathbb{R}^{L\times D}$, it first undergoes a bottleneck 1D convolution to compress its spatial dimension from $D$ to $G$, where compression factor $G\ll D$. Multi-scale temporal information is then obtained and merged via FCTF across four different temporal scales to generate fine-to-coarse temporal feature $f_{FC}$. After that, the $f_{FC}$ is expanded with a linear layer and split into two information flows. The first flow passes through a 1D convolutional layer, followed by a SiLU activation function [34] before entering the selective SSM. The output of the selective SSM is then gated by the $z$ generated from the other flow. Finally, we use another bottleneck 1D convolutional layer to restore the spatial dimension to $\frac{D}{2}$ and get the output of the BMSS block $y\in\mathbb{R}^{L\times\frac{D}{2}}$. Formally, we present the operations of the BMSS block in Algorithm 1. The BMSS block inherits the linear complexity of the selective SSM while retaining a local-to-global receptive field, which is promising for video-based temporal error detection.
Algorithm 1 Bottleneck Multi-scale State Space (BMSS) Block
[…]
2:Output: $y:(L,D/2)$.
3:/* Bottleneck: Compress $x$ along spatial dimension */
4:$f_{C}:(L,G)\leftarrow\text{1DConv}^{G}(x)$
5:/* Fine-to-Coarse Temporal Fusion (FCTF): Obtain
6:    multi-scale temporal information across four scales*/
7:$f_{T1}:(L,E)\leftarrow\text{D1DConv}^{E}(f_{C})$ with 2 dilation rate
8:$f_{T2}:(L,E)\leftarrow\text{D1DConv}^{E}(f_{T}1)$ with 4 dilation rate
9:$f_{T3}:(L,E)\leftarrow\text{D1DConv}^{E}(f_{T}2)$ with 8 dilation rate
10:$f_{FC}:(L,G)\leftarrow\text{1DConv}^{G}\left(\text{Cat}(f_{C},f_{T1},f_{T2},f_{T3})\right)$
11:/* Information split */
12:$x:(L,2G)\leftarrow\text{Linear}^{2G}(f_{FC})$
13:$x,z:(L,G)\leftarrow\text{Split}(x)$
[…]
15:/* Selective SSM*/
16:$\mathbf{A}:(G,N)\leftarrow\text{Parameter}(\mathbf{A})$
17:$\mathbf{B}:(L,N)\leftarrow\text{Linear}_{\mathbf{B}}^{N}(x)$
18:$\mathbf{C}:(L,N)\leftarrow\text{Linear}_{\mathbf{C}}^{N}(x)$
[…]
20:                  $\left.\left.\text{Broadcast}^{G}(\text{Linear}_{\Delta}^{1}(x))\right)\right)$
21:$\overline{\mathbf{A}},\overline{\mathbf{B}}:(L,G,N)\leftarrow\text{Discretization}(\Delta,\overline{\mathbf{A}},\overline{\mathbf{B}})$
22:$x:(L,G)\leftarrow\text{SSM}(\overline{\mathbf{A}},\overline{\mathbf{B}},\mathbf{C})(x)$
23:/* Get gated $y$/
24:$y:(L,G)\leftarrow\text{Linear}^{G}(x\otimes\text{SiLU}(z))$
25:/* Bottleneck: Restore $y$ along spatial dimension */
26:$y:(L,D/2)\leftarrow\text{1DConv}^{D/2}(y)$
27:return $y:(L,D/2)$
II-B4 Fine-to-Coarse Temporal Fusion (FCTF)
To tackle varying durations of error and restricted receptive fields of convolutional layers, we propose the FCTF to capture and merge multi-scale temporal information. FCTF stacks three dilated 1D convolution layers with dilation rates of 2, 4, and 8 from top to bottom, enabling the fine-to-coarse temporal feature extraction scale by scale. Besides, stacking dilated convolution increases the receptive field without the need to increase the parameter number by increasing the kernel size. The receptive field expands exponentially as the number of stacked layers increases. Therefore, with a few parameters, we achieve a significantly large receptive field in the temporal dimension, which mitigates model overfitting and facilitates efficient error detection. The receptive field at each layer, with a kernel size of 3, can be calculated with the following formula:
$\text{ReceptiveField}(l)=2^{l+2}-1$ (5)
where $l\in[1,3]$ represents the layer number.
Specifically, given the compressed feature $f_{C}\in\mathbb{R}^{L\times G}$, which is obtained via the first bottleneck 1D convolution layer. It first goes through a 1D convolution layer with 2 dilated rate, which extracts a temporal feature of a receptive field size of 7, denoted as $f_{T1}$. Similarly, the $f_{T1}$ undergoes further refinement through dilated 1D convolution layers with 4 and 8 dilation rates, thereby acquiring temporal features $f_{T2}$ and $f_{T3}$ across larger and coarser time spans. These temporal features of varying granularity are concatenated with the original scale $f_{C}$ along the spatial dimension, and ultimately, a 1D convolution with 1 kernel size fuse them into a fine-to-coarse temporal feature $f_{FC}\in\mathbb{R}^{L\times G}$.
III Experiments
III-A Experimental Settings
All experiments are implemented in PyTorch on a GTX 4090Ti GPU. We use AdamW optimizer with initial learning rate of 1e-4 and maximum training epoch set at 200. We sampled SAR-RARP videos at 5Hz to get 40 videos with 46,586 normal frames and 18,540 error frames as the training set, and 8 videos with 7549 normal frames and 2096 error frames as the testing set. We use DINOv2 with fixed weights pretrained on ImageNet [33] as the visual extractor to generate spatial embedding sequences $\mathcal{E}$ from the video with $L$ frames as the model input. The compression factor $G$ in the BMSS block is set as 64. The channel $E$ of dilated 1D convolutional layers in FCTF is fixed as $\frac{D}{8}$. Following  [27], the state dimension $N$ of selective SSM is set to $16$. We the standard and widely used evaluation metrics: area under the curve (AUC) and average precision (AP) for the SED task. We report the mean and standard deviation from three runs with different random seeds.
III-B Comparison with the State-of-the-Art
We compare the proposed SEDMamba with state-of-the-art convolutional and Transformer-based video analysis methods: TeCNO [35], MS-TCN [36], MS-TCN++ [37], ASFormer [24], and selective SSM-based methods: Mamba [27] and Vim [28]. All methods are implemented based on their released code and original literature, and fine-tuned to fit our error detection task to provide a basis for a fair comparison.
Table II shows quantitative results of all methods. The proposed SEDMamba consistently achieves superior performance with an increase of 1.82-13.06% in AUC and 3.8-18.86% in AP, compared to other methods. Notably, all selective SSM-based methods surpass the Transformer-based method ASFormer, demonstrating the superiority of selective SSM in the SED task. Fig. 4 illustrates error detection results from all test videos. The second-best method Mamba tends to output smoother and more ambiguous prediction curves, which impedes its sensitivity to errors. It also produces lower confidence outputs within error frames. Thanks to the fine-to-coarse temporal information provided by FCTF for selective SSM, SEDMamba is able to identify suddenly occurring errors with short duration (highlighted in red boxes) and consistently produces high confidence within error areas in most cases (highlighted in black boxes). In most normal frames, SEDMamba has comparable or lower error probabilities to Mamba, with only a small portion showing higher error probabilities. Note that both AUC and AP comprehensively consider all classes, i.e. error and normal, indicating SEDMamba has superior overall performance.
TABLE II: Quantitative results of SEDMamba and other methods. Higher AUC and AP mean better performance. Top two results are highlighted in bold and underlined.
Methods AUC (%) $\uparrow$ AP (%) $\uparrow$
TeCNO [35] 58.14$\pm$0.33 26.01$\pm$0.32
MS-TCN [36] 58.82$\pm$0.16 29.98$\pm$3.06
MS-TCN++ [37] 58.57$\pm$0.43 30.35$\pm$1.35
ASFormer [24] 68.76$\pm$0.93 36.79$\pm$2.35
Mamba [27] 69.38$\pm$0.26 41.07$\pm$0.95
Vim [28] 69.06$\pm$1.46 40.63$\pm$2.30
SEDMamba (ours) 71.20$\pm$0.26 44.87$\pm$1.52
Figure 4: Visualization results of the proposed SEDMamba (green line) and the second-best method Mamba (blue line) on all 8 complete test videos. Each row represents one whole test video. The X-axis is the frame index while the Y-axis is the error probability output by the models. The red translucent background indicates error frames, while the white background indicates normal frames, i.e. ground truth of each frame. Different coloured bars on the X-axis represent different gestures. Red and black boxes are used to highlight significant results.
III-C Computational Efficiency
Given a video spatial embedding sequence $\mathcal{E}\in\mathcal{R}^{L\times D}$, the computational complexity of a single-head self-attention Transformer and a selective SSM in SEDMamba are $O(L^{2}D+LD^{2})$ and $O(LDN)$, respectively. Computationally, self-attention is quadratic to sequence length $L$, while selective SSM is linear to $L$. This computational efficiency makes SEDMamba scalable for gigapixel applications with large video lengths. Besides, as shown in Table II and Table III, SEDMamba achieves the highest performance yet maintains the lowest parameters and FLOPs. Specifically, compared to other selective SSM-based methods, Mamba and Vim, our SEDMamba obtains 58.23% and 75.79% model parameter reductions, and 51.79% and 70.90% FLOPs reductions, respectively. Moverover, the parameters and FLOPs of our SEDMamba are only 27.38% and 9.07% of those of the Transformer-based method ASFormer, respectively, yet it achieves a performance gain of 2.44% in AUC and 8.08% in AP. Compared to methods based on 1D and dilated convolution, i.e. TeCNO, MS-TCN and MS-TCN++, our method gets far more performance improvement with a smaller complexity. This is attributed to the bottleneck construction of BMSS blocks, which compresses the spatial dimension at the input and restores it at the output for computational saving, as well as the FCTF therein BMSS, which provides fine-to-coarse temporal features for selective SSM to model long-term dependencies effectively.
TABLE III: Computational complexity of SEDMamba and other methods in terms of model parameters (Params) and floating point operations (FLOPs). Lower Params and FLOPs mean lower model computational complexity. K = ×$\times 10^{3}$, M = ×$\times 10^{6}$.
Methods Params (K) $\downarrow$ FLOPs (M) $\downarrow$
TeCNO [35] 328.51 45.17
MS-TCN [36] 725.19 71.98
MS-TCN++ [37] 922.63 91.64
ASFormer [24] 1059.27 295.37
Mamba [27] 694.27 55.55
Vim [28] 1197.83 92.03
SEDMamba (ours) 290.03 26.78
III-D Ablation Study
III-D1 Effectiveness of Key Components in BMSS Block
We study the effect of bottleneck mechanism and FCFT in BMSS blocks. As shown in Table IV, whereby removing each component leads to performance degradation. The bottleneck mechanism yields performance gains of +1.64% in AUC and +3.69% in AP, while model parameters and FLOPs are significantly decreased by 82.98% and 78.83%, which guards against overfitting and enhances model efficiency. Besides, such a bottleneck mechanism compresses features on the spatial dimension, which can mitigate the impact of redundant spatial noise on long-term dependency modelling. Introducing FCTF to bottleneck mechanism leads to a slight increase in computational complexity, which is acceptable given the substantial performance gains of +1.58% in AUC and +3.58% in AP. This underscores the importance of providing multi-scale temporal information for selective SSM in the surgical error detection task.
TABLE IV: Ablation results for bottleneck mechanism and FCFT in BMSS blocks.
Bottleneck FCFT Params (K) $\downarrow$ FLOPs (M) $\downarrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$
1569.79 107.05 67.98$\pm$1.68 37.60$\pm$1.80
267.11 22.66 69.62$\pm$0.65 41.29$\pm$0.84
290.03 26.78 71.20$\pm$0.26 44.87$\pm$1.52
TABLE V: Comparison of different numbers of BMSS blocks.
Blocks Params (K) $\downarrow$ FLOPs (M) $\downarrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$
1 137.35 19.29 69.96$\pm$1.29 41.47$\pm$2.29
2 225.69 30.09 70.44$\pm$0.49 43.23$\pm$0.78
3 290.03 26.78 71.20$\pm$0.26 44.87$\pm$1.52
4 342.34 41.16 66.44$\pm$1.44 38.01$\pm$0.80
5 388.64 45.60 66.21$\pm$2.76 36.32$\pm$2.73
TABLE VI: Comparison of different compression factor $G$ in BMSS blocks.
G Params (K) $\downarrow$ FLOPs (M) $\downarrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$
16 54.61 7.45 67.60$\pm$1.74 37.57$\pm$5.02
32 120.65 12.99 68.32$\pm$0.85 39.80$\pm$2.46
64 290.03 26.78 71.20$\pm$0.26 44.87$\pm$1.52
128 777.98 65.12 71.84$\pm$0.87 45.17$\pm$1.47
III-D2 Number of BMSS Blocks
As shown in Table V, the depth of our SEDMamba can be tailored by hierarchically stacking BMSS blocks. Remarkably, even with a single BMSS block, our method achieves an AUC of 69.96% and an AP of 41.47%, surpassing the second-best method Mamba, while our method’s parameters and FLOPs are only 19.78% and 34.73% of Mamba’s. This further verifies the effectiveness of the proposed BMSS block. Performance continues to improve with more BMSS blocks but declines when the model becomes too deep, such as with five BMSS blocks. The potential reason behind this decrease could be the gradient vanishing/exploding issue commonly found in deep models. Introducing residual connections between BMSS blocks might be helpful.
III-D3 Compression Factor in Bottleneck Mechanism
The compression factor $G$ of the bottleneck mechanism in the BMSS block determines the spatial dimension of the input for the FCTF and selective SSM. As shown in Table VI, it is evident that altering $G$ significantly affects the computational complexity of the model. Considering the trade-off between performance and computational complexity, we set $G$ to 64.
IV Conclusion
This paper presents the first frame-level error annotations in an in-vivo surgical dataset. More than 970000 frames from 48 videos in SAR-RAP50 are annotated frame by frame to provide temporal boundaries of errors. Based on this, we propose SEDMamba, a hierarchical surgical error detection model based on selective SSM with bottleneck and fine-to-coarse temporal fusion, which demonstrates improvements in error detection performance with high computational efficiency. Experimental results show that the proposed method outperforms six state-of-the-art competitions with 1.82-13.06% AUC and 3.8-18.86% AP performance improvements and better qualitative results. Exploiting a very efficient design SEDMamba requires between 11.71-75.79% less model parameters leading to a 40.71-90.93% FLOPs reduction. Future work will focus on detecting error types semantically with multi-label algorithms.